### v1

# Unlocking Personalized Knowledge in Federated Large Language Model: The Power of Mixture of Experts

Abstract
The Mixture of Experts (MoE) architecture has emerged as a prominent strategy for scaling large language models (LLMs), effectively leveraging sparse activation and facilitating task-specific personalization. However, current federated learning (FL) approaches are primarily designed for dense models, making them unable to directly exploit the sparsity inherent in MoE architectures. Treating MoE models as dense networks in federated scenarios results in excessive communication overhead and computational costs, undermining the potential for personalized knowledge sharing. To address these challenges, we propose FLEx (Federated LLMs with Personalized Experts), a novel federated learning framework explicitly tailored for MoE-based LLMs. FLEx efficiently personalizes by pruning the global MoE model to keep only one expert per client, and employs an adaptive gating mechanism to reintegrate these personalized experts into the pre-trained MoE layers, ensuring the original backbone architecture remains unchanged. These personalized experts are trained with local data and stored locally on each client, while the shared modules are aggregated globally. Extensive evaluations on diverse instruction-based datasets under non-IID conditions consistently demonstrate that FLEx outperforms existing federated baselines. Our code is available at https://anonymous.4open.science/r/FLEx-8F12.
1 Introduction
Large language models (LLMs) have seen a remarkable increase in both scale and capability in recent years, largely due to the rise of the Mixture-of-Experts (MoE) paradigm jacobsAdaptiveMixturesLocal1991 ; fedusSwitchTransformersScaling2022 ; jiangMixtralExperts2024 . By activating only a small subset of expert sub-networks per token, MoE architectures separate the number of parameters from the computational cost, enabling state-of-the-art performance at significantly lower inference overhead compared to dense Transformers of similar size. Recent advances deepseek-aiDeepSeekV3TechnicalReport2025 ; qwen2.5 highlight that MoEs are not just an efficient compression strategy, but a powerful means of task-specific personalization, where experts naturally specialize in different linguistic styles, domains, or user preferences.
Meanwhile, federated learning (FL) has become a compelling framework for training and adapting LLMs on privacy-sensitive data that remains on users’ devices zhangBuildingFederatedgptFederated2024 ; yeOpenFedLLMTrainingLarge2024 ; longDualpersonalizingAdapterFederated2024 ; wangFLoRAFederatedFineTuning2024 . However, nearly all existing FL algorithms for LLMs are designed for dense model architectures. Applying them directly to MoE models requires each client to transmit all experts, even those unrelated to its local data, resulting in high communication costs and diminished personalization. Early attempts to bridge MoE and FL zouFedMoEEfficientFederated2025 ; meiFedMoEPersonalizedFederated2024 ; luoMixtureExpertsMade2025 were developed for traditional models rather than trillion token pretrained LLMs. While they prove that federated MoE training is feasible, they still require each client to manage and transmit multiple experts, including those that are not activated by local data, resulting in unnecessary communication overhead, which makes these methods economically impractical and leaves a significant efficiency gap in federated training of large-scale pretrained LLMs.
Deploying MoE-based LLMs in a federated setting introduces two main challenges. First, the communication and computation costs that arise from the sparse structure of MoE models are high, especially under limited bandwidth and constrained local device resources. Second, existing methods primarily focus on personalized experts trained from local data while neglecting global information, failing to achieve an effective balance between personalization and global knowledge within MoE-based LLMs.
To address these challenges, we propose FLEx (Federated LLMs with Personalized Experts), a simple framework that combines the sparsity of Mixture-of-Experts models with the benefits of federated personalization. In FLEx, each client selects and fine-tunes only one expert from the global MoE model, tailored to its local data. During fine-tuning, a lightweight gating module integrates the personalized expert’s output into the original MoE layer, allowing the model to leverage expert specialization without altering the backbone architecture. In FLEx, clients locally update their chosen expert and specific parameters in the shared module. Only the shared module updates are then communicated for aggregation into the global model, ensuring shared improvements, while the expert’s modifications remain local to the client for personalization. Our main contributions are as follows:
  • We propose FLEx, a novel federated learning framework explicitly tailored for MoE-based large language models. FLEx addresses the lack of federated learning approaches specifically designed for sparse MoE architectures, effectively balancing global knowledge sharing and local personalization, thereby significantly reducing communication and computational costs compared to traditional FL methods.
  • FLEx efficiently personalizes by pruning the MoE model to retain a single expert per client and using an adaptive gating mechanism to reintegrate personalized experts into the pre-trained MoE layers without altering the original backbone architecture. These personalized experts are trained using local data and stored locally on each client, while the shared modules are aggregated globally. This design maximally preserves pre-trained knowledge, leverages personalized experts to enhance client-specific capabilities, benefits from shared global knowledge aggregation, and significantly reduces communication overhead.
  • Extensive evaluations conducted on multiple challenging instruction-based datasets, including Databricks-dolly-15k, Alpaca-GPT4, Finance-Alpaca, and MedAlpaca, under non-IID scenarios demonstrate that FLEx consistently outperforms existing federated learning baselines across diverse tasks.
2 Related Work
2.1 Federated Learning on Large Language Models
Federated Learning (FL) has gained significant attention as a promising method for training Large Language Models (LLMs), addressing challenges such as decentralized data management and privacy concerns zhangBuildingFederatedgptFederated2024 ; yeOpenFedLLMTrainingLarge2024 ; wangFLoRAFederatedFineTuning2024 . Recent surveys yaoFederatedLargeLanguage2024 ; yeOpenFedLLMTrainingLarge2024 have identified important intersections between FL and LLMs, particularly in areas such as efficient foundation model training, federated fine-tuning techniques, and the collaborative potential of FL in advancing LLM development. Notable contributions in this domain include FedIT zhangBuildingFederatedgptFederated2024 , which demonstrated the utility of FL in instruction tuning tasks for LLMs. Further innovations involve personalized LLM training methods like DualLoRA longDualpersonalizingAdapterFederated2024 and FDLoRA qiFDLoRAPersonalizedFederated2024 , which combine federated learning with personalized adapter methods. Other approaches such as FRLoRA yanFederatedResidualLowRank , FLoRA wangFLoRAFederatedFineTuning2024 , and FlexLoRA baiFederatedFinetuningLarge introduce new aggregation strategies designed to optimize low-rank fine-tuning and allow more flexible model updates in federated settings. Additionally, selective aggregation methods, like those presented in guoSelectiveAggregationLowRank2025 , further improve federated LLM training by refining the aggregation process.
2.2 Mixture of Experts
Mixture-of-Experts (MoE) jacobsAdaptiveMixturesLocal1991 ; fedusSwitchTransformersScaling2022 models activate only a subset of expert layers during inference or training, which significantly reduces computational costs while maintaining high performance. In recent, MoE architectures have gained attention for their scalability and efficiency in large language models jiangMixtralExperts2024 ; deepseek-aiDeepSeekV3TechnicalReport2025 ; qwen2.5 .
However, existing federated MoE approaches zhanFedMoEDAFederatedMixture2024 ; luoMixtureExpertsMade2025 are primarily designed for dense model architectures. Applying these methods directly to MoE models can result in substantial computational and communication overhead. To address this limitation, we propose FLEx, a novel framework specifically designed for federated training of MoE-based LLMs.
2.3 Expert Pruning for LLMs
Model pruning is a widely used technique to improve communication and computation efficiency in deep learning. In the context of MoE-based LLMs, recent pruning methods, such as Wanda sunSimpleEffectivePruning2024 and SparseGPT frantarSparsegptMassiveLanguage2023 , offer effective strategies to reduce the number of model parameters. Additionally, recent studies luNotAllExperts2024 ; liuEfficientExpertPruning2024 have demonstrated that many experts are either unimportant or redundant during inference on specific benchmarks. However, in federated learning with MoE-based LLMs, directly pruning certain experts may result in the loss of essential foundational knowledge. To address this, we propose pruning personalized side experts on local data. Furthermore, we integrate the pruned experts back into the original pretrained model to achieve an optimal balance between local adaptation and global generalization.
Figure 1: Overview of FLEx framework. The FLEx framework begins by pruning personalized experts for each client using local data. The next step involves injecting personalized knowledge into the MoE layer via a gating mechanism, striking a balance between global knowledge sharing and local adaptation.
3 Preliminary
Federated Learning Setup
We consider a federated learning (FL) setup with $K$ clients, each holding a private dataset $\mathcal{D}_{k}$. In this setup, each client independently trains a model on its local data, while the central server periodically aggregates the updates to improve the global model. A widely adopted approach, the Federated Averaging (FedAvg) algorithm, aggregates the global parameters by computing a weighted average of the local parameters:
$W^{t+1}=\sum_{i=1}^{N}\frac{n_{i}}{n}W_{i}^{t+1},$ (1)
where $n=\sum_{i=1}^{N}n_{i}$ is the total number of training samples across all clients, and $W^{t+1}_{i}$ and $W^{t+1}$ are weight of $i$-th client and server in $t+1$-th round. This approach enables the global model to leverage distributed data while maintaining privacy.
Mixture of Experts in Large Language Models
The Mixture of Experts (MoE) architecture introduces a sparse activation mechanism, where only a subset of experts is activated for each input. This design enables large-scale models to scale efficiently while reducing computational costs. In modern LLMs, the backbone consists of transformer layers, each comprising a self-attention layer and a feedforward network (FFN) layer:
[…]
$h_{t}^{l}$ $=\text{FFN}\left(u_{t}^{l}\right)+u_{t}^{l},$ (3)
where $u_{1:T}^{l}$ is hidden states of all tokens after the $l$-th attention module and $h_{t}^{l}$ is the output hidden state of the $t$-th token after the $l$-th Transformer block. In MoE-based models, the FFN layer is replaced by a MoE layer, which consists of multiple experts. This substitution yields:
$h_{t}^{l}=\sum_{i=1}^{N}\left(g_{i,t}\text{FFN}_{i}\left(u_{t}^{l}\right)\right)+u_{t}^{l},$ (4)
where $N$ denotes the total number of experts, and $g_{i,t}$ is the gating function that selects which experts are activated for each token $t$. The gating function $g_{i,t}$ is defined as:
$g_{i,t}=\begin{cases}s_{i,t},&\text{if }s_{i,t}\in\text{Topk}\left(\{s_{j,t}|1\leq j\leq N\},K\right)\\ 0,&\text{otherwise},\end{cases}$ (5)
where $s_{i,t}$ is the relevance score for expert $i$ for token $t$, computed using the softmax function:
$s_{i,t}=\text{Softmax}_{i}\left(\text{router}(u_{t})\right).$ (6)
Here, the router function $\text{router}(u_{t})$ computes the raw scores for each expert based on the token’s representation $u_{t}$, and the softmax function $\text{Softmax}_{i}$ normalizes these scores to determine the relevance of each expert. The gating mechanism activates only the top $K$ experts for each token $t$, which significantly reduces computational costs. However, if we treat MoE-based LLMs as dense models and aggregate the entire model, the communication cost would grow significantly, making it impractical. To address this, we propose FLEx and unlock the power for experts in LLMs.
4 Unlock the Power of Experts: FLEx
Federated learning (FL) enables distributed model training across multiple clients, each holding a private dataset. Mixture of Experts (MoE) models employ sparse activation, activating only a small subset of experts for each input. This sparsity makes it impractical to aggregate the entire model across all clients due to significant communication overhead. In this paper, we propose FLEx, a federated learning approach that optimizes MoE models by pruning the MoE layer to retain a personalized expert for each client. FLEx reduces communication costs, maintains high performance, and preserves knowledge from the original pre-trained MoE model. By integrating personalized experts into each client’s model, FLEx strikes a balance between global model benefits and local data needs. The framework of FLEx is shown in Figure 1.
4.1 Selective Aggregation in FLEx
In MoE-based LLMs, communication efficiency is crucial. FedAvg, the standard federated training algorithm, faces challenges when aggregating both dense attention layers and sparse MoE layers. While MoE layers activate only a subset of experts for each input, the attention layers in LLMs are dense and require substantial resources for aggregation in a federated setup. To address this, we propose a hybrid approach. The attention layers remain shared and aggregated globally, while the MoE layers are pruned locally. Instead of aggregating all MoE parameters, we fix the MoE layer’s weights and prune it to retain a personalized expert for each client.
4.2 Personalized Experts via MoE Pruning
To further enhance personalization and computational efficiency in federated learning, inspired by luNotAllExperts2024 , we introduce a pruning strategy based on MoE layers. This approach allows each client to select the most suitable subset of experts for their local data, maximizing personalization while avoiding the redundancy and resource consumption that may arise during global aggregation. Next, we will detail the specific steps and optimization strategies involved in this process.
Each client $i$ computes the output of all experts for a given input $x\in\mathcal{D}_{i}$. Let $F^{(l)}(x)$ denote the output of the MoE layer, and let $S\subset\{1,\dots,K\}$ represent the subset of experts retained after pruning. The pruned output at layer $l$ is given by $F^{\prime(l)}(x,S)$.
For personalization, each client selects the subset of experts that best approximates the original MoE output for their local dataset at each layer by minimizing the reconstruction loss, as shown in Step 1 of Figure 1. This optimization problem for a given layer $l$ is formulated as:
$S_{i}^{(l)}=\arg\min_{S\subset\{1,\dots,K\},|S|=n}\left(\frac{1}{|\mathcal{D}_{i}|}\sum_{x\in\mathcal{D}_{i}}\|F_{S}^{(l)}(x)-F^{(l)}(x)\|_{F}\right).$
To simplify the problem, we reduce the search space by setting $n=1$, choosing only a single expert per client. This reduces the computational complexity while ensuring effective personalization. Through this optimization problem, we greedyly select the personalized side experts at each layer.
4.3 Personalized Layer Integration
Once the personalized expert is selected, it is integrated into the global model by combining it with the global MoE layer, as shown in Step 2 of Figure 1. During the federated aggregation process, the LoRA adapter for the attention layer is aggregated globally, while the LoRA adapter for the personalized expert remains on the client side. This approach ensures that each client benefits from both the global model’s knowledge and the personalized expert’s adaptation to local data. Then the forward equation of MoE layer (4)
$h_{t}^{l}$ $=\sum_{i=1}^{N}\left(g_{i,t}\text{FFN}_{i}\left(u_{t}^{l}\right)\right)+g_{e}^{t}\text{FFN}_{e}(u_{t}^{l})+u_{t}^{l},$ (7)
$g_{e}^{t}$ $=\text{sigmoid}(\text{router}_{e}(u_{t}^{l})),$ (8)
where $\text{router}_{e}$ and $\text{FFN}_{e}$ denote the personalized router and personalized side expert generated by pruned experts, respectively. Here, we keep the original FFN layers fixed to ensure that our method preserves as much knowledge as possible for trillion-token pretrained MoE-based LLMs.
Remark.
Our method retains the pre-trained weights in the MoE layer while adding a personalized side expert. This design maximally preserves pre-trained knowledge, leverages personalized experts to enhance client-specific capabilities, benefits from shared global knowledge aggregation, and significantly reduces communication overhead.
5 Experiments
In this section, we will use substantial experiments to demonstrate the effectiveness of FLEx.
5.1 Experimental Setup
Datasets and Data Partition.
To simulate real-world scenarios, we utilize the Databricks-dolly-15k [^1] https://huggingface.co/datasets/databricks/databricks-dolly-15k dataset, an open-source collection of instruction follow-up records generated by numerous Databricks employees. This dataset covers a variety of behavioral categories as described in the InstructGPT paper ouyangTrainingLanguageModels2022 and includes eight subtasks. We select four subtasks—classification, closed QA, information extraction, and summarization—to create a non-IID scenario. These tasks have fixed responses, unlike open QA or creative writing tasks. Therefore, ROUGE-L serves as an effective evaluation metric to assess the instruction followability capability of large language models.
Implementation.
We utilize the OpenFedLLM framework yeOpenFedLLMTrainingLarge2024 to evaluate federated learning methods, including FedAvg mcmahanCommunicationefficientLearningDeep2017 , FedProx liFederatedOptimizationHeterogeneous2020 , SCAFFOLD karimireddyScaffoldStochasticControlled2020 , FedAvgM hsuMeasuringEffectsNonIdentical2019 , FedAdagrad reddiAdaptiveFederatedOptimization2021 , FedYogi reddiAdaptiveFederatedOptimization2021 and FedAdam reddiAdaptiveFederatedOptimization2021 . Each client trains on 20 samples per round for 100 rounds. We select the same round checkpoint for each client to compare ROUGE-L scores. The Qwen1.5-MoE-A2.7B model [^2] https://huggingface.co/Qwen/Qwen1.5-MoE-A2.7B, configured with bfloat16, serves as the base model and is trained on NVIDIA H20 GPUs with 96GB of memory. We apply LoRA to the attention layer with a rank of 32 and an alpha of 64. The Adam optimizer is used to train the LoRA adapter with a batch size of 4, using a consistent learning rate of 4e-5. The maximum input sequence length is set to 2048, and we provide the Alpaca template as a prompt.
Table 1: Rouge-L performance comparison of different federated learning strategies, including our proposed method, on the Databricks-dolly-15k dataset under a non-IID scenario (one subtask per client). The results are reported for four distinct instruction-based tasks: classification, closed QA, information extraction, and summarization, using the Rouge-L metric to evaluate instruction followability. The best result over federated method is bolded and the second-best result is underlined. CLF: Classification, CQA: Closed Question Answering, IE: Information Extraction, Summ: Summarization
Databricks-dolly-15k MMLU
Method CLF CQA IE Summ Avg Avg
Local Training 51.90 34.34 40.93 40.37 41.88 44.32
MoE + FedAvg 51.39 34.52 37.23 41.54 41.17 43.91
MoE + FedAvgM 51.69 36.72 38.54 42.54 42.37 40.13
MoE + FedAdam 49.77 35.88 37.45 41.11 41.05 42.57
MoE + FedAdagrad 50.09 36.38 38.71 42.24 41.85 47.06
MoE + SCAFFOLD 51.20 34.45 37.93 41.75 41.33 45.01
MoE + FedProx 51.55 34.71 35.07 41.71 40.76 44.79
MoE + FedYogi 49.73 36.67 36.27 41.36 41.00 44.05
FLEx (Ours) 52.10 38.23 40.13 42.09 43.13 49.74
5.2 Performance under Pathological Non-IID Distribution
We first consider pathological data heterogeneity, where we assign one subtask per client, with a total of four clients. As shown in Table 1, we compared our methods with baseline methods. From this table, we can observe that FLEx (Ours) has the highest average performance (43.13), outperforming all other methods. From an averaging perspective, MoE+FedAvgM (42.37) is the next closest competitor. Also, performance seems to vary by task, except summarization task, FLEx achieves the highest score. From this table, we can observe that the proposed method consistently leads in average performance, indicating its effectiveness across various tasks. Besides, it can be observed that due to the powerful pretrained model, simply apply averaging as traditional federated paradigm will easily cannot outperform simple local update and only our method and MoE+FedAvgM can ourperform local update.
We also evaluated the performance of each client’s checkpoint on the MMLU-STEM benchmark wangMMLUPRO to assess the extent of knowledge forgetting after personalization. Specifically, each client’s model was tested individually, and the average performance was calculated. The results indicate that FLEx (Ours) preserves the most general knowledge while maintaining strong performance in personalization. In contrast, MoE+FedAvgM, which showed competitive results in terms of personalization, demonstrated the greatest loss of general knowledge, highlighting its tendency to forget important shared knowledge during the personalization process. This further demonstrates the effectiveness of our FLEx method, which strikes a better balance between general knowledge retention and personalization.
Table 2: Rouge-L performance comparison of different federated learning strategies, including our proposed method, on the Alpaca-gpt4, Finance-Alpaca and MedAlpaca dataset under a non-IID scenario (one subtask per client). The results are using the Rouge-L metric to evaluate instruction followability. The best result over federated method is bolded and the second-best result is underlined.
Method Alpaca-gpt4 Finance-Alpaca MedAlpaca Avg
Local Training 31.31 28.64 30.93 30.29
MoE + FedAvg 30.11 28.17 30.06 29.44
MoE + FedAvgM 28.90 27.18 28.78 28.28
MoE + FedAdam 29.23 28.70 31.18 27.70
MoE + FedAdagrad 30.24 28.79 30.60 29.87
MoE + SCAFFOLD 29.96 28.38 28.92 29.08
MoE + FedProx 30.05 28.60 30.01 29.55
MoE + FedYogi 29.73 28.42 30.51 29.55
FLEx (Ours) 31.54 29.89 31.31 30.91
Table 2 presents the ROUGE-L performance comparison of different federated learning strategies, including our proposed method, FLEx, on datasets that simulate real-world user scenarios: Alpaca-gpt4, Finance-Alpaca, and MedAlpaca. The experiment was conducted in a highly non-IID setting, where each client was assigned a domain-specific dataset.
FLEx outperforms all other methods across all datasets, achieving the highest ROUGE-L scores: 31.54 for Alpaca-gpt4, 29.89 for Finance-Alpaca, and 31.31 for MedAlpaca, with an average score of 30.91. This demonstrates its robustness in handling diverse and specialized data. The second-best performing method, MoE + FedAdagrad, shows an average score of 29.87, but its performance is inconsistent. While it ranks second on MedAlpaca (31.18), its performance on Alpaca-gpt4 and Finance-Alpaca is lower, highlighting its instability. Interestingly, federated averaging methods (e.g., MoE + FedAvg) and other federated optimizers (MoE + FedAvgM, MoE + FedAdam) fail to outperform the local update method (Local), which achieves an average score of 30.29. This emphasizes that traditional federated learning approaches struggle to effectively aggregate models in highly non-IID scenarios, sometimes even underperforming compared to isolated local training.
These results reinforce the rationale behind our choice of domain-specific datasets like Finance-Alpaca and MedAlpaca, which simulate the specialized needs of real-world users. FLEx excels in such complex, heterogeneous data environments by capturing personalized knowledge and effectively leveraging domain-specific features, outperforming both baseline methods and local training.
5.3 Performance under Dirichlet Non-IID Distribution
Compared to previous experiments on pathological non-IID scenerios, we introduce dirichlet non-IID distributions into our experiments as shown in Table 3. In federated learning (FL), the Dirichlet non-IID distribution refers to the scenario where data is non-identically and independently distributed across clients, with each client’s data distribution being generated from a Dirichlet distribution, typically used to model the unevenness of data distribution in real-world scenarios. Here, we choose $\alpha=1.0$ as our dirichlet distribution parameter.
We introduced the Dirichlet non-IID distribution to further evaluate the robustness of our proposed FLEx method in a more realistic and challenging scenario. The Dirichlet distribution, with its ability to model the unevenness of data distribution across clients, is commonly used in federated learning to simulate real-world data heterogeneity. By increasing the number of clients to 10 in this experiment, we aimed to simulate a more complex and distributed environment, which is closer to practical deployment scenarios in federated learning.
Table 3: ROUGE-L performance comparison of different federated learning strategies, including our proposed method, on the Databricks-dolly-15k dataset under a Dirichlet distribution with 10 clients, aggregating 4 clients per communication round. The results are reported for four distinct instruction-based tasks: classification, closed QA, information extraction, and summarization, using the ROUGE-L metric to evaluate instruction following ability.
Databricks-dolly-15k
Method alpha=0.1 alpha=1.0
Local Training 35.18 36.88
MoE + FedAvg 35.29 37.51
MoE + FedAvgM 37.41 35.94
MoE + FedAdam 34.69 35.79
MoE + FedAdagrad 36.84 36.84
MoE + SCAFFOLD 34.08 36.91
MoE + FedProx 35.05 36.68
MoE + FedYogi 35.26 36.41
FLEx (Ours) 37.30 37.54
To ensure fairness and consistency, we selected a fixed communication round and evaluated the performance using the aggregated checkpoints of four clients per round, chosen based on their data distributions. For each task, the client whose data distribution most closely matches the task’s focus was selected. For example, if a client’s dataset contained the majority of samples for a specific task, the evaluation for that task was conducted based on the corresponding client’s model. This approach ensures that the task-specific performance is reflective of the data the clients are primarily working with, making the evaluation more aligned with the real-world distribution of data. When analyzing the results for $\alpha=0.1$, FLEx (37.30) ranks closely behind the top method, MoE + FedAvgM (37.41). It is important to note that as a personalized federated learning approach, FLEx is evaluated on each client’s personalized checkpoint, rather than a single global model. This distinction influences its performance relative to methods optimized solely for global averages. The relatively close performance values observed for $\alpha=1.0$ in this experiment are characteristic of the Dirichlet distribution; higher $\alpha$ values in this distribution lead to more balanced data heterogeneity across clients, which tends to reduce the performance gap between different FL methods compared to pathological non-IID scenarios. The key takeaway is that FLEx remains competitive even in scenarios with less extreme data heterogeneity. Its ability to effectively balance personalization with adaptation to a broader data distribution underscores its robustness and adaptability to diverse client needs, standing out from methods focused solely on global performance.
5.4 Performance under LLM-as-a-Judge Evaluation
Vicuna
Method Helpfulness Harmlessness
Base Model 5.145 6.319
Local Training 6.026 7.096
MoE + Fedavg 6.159 7.661
MoE + FedAvgM 6.043 7.527
MoE + FedAdam 6.267 7.577
MoE + FedAdagrad 6.314 7.780
MoE + SCAFFOLD 5.908 7.248
MoE + FedProx 5.949 7.475
MoE + FedYogi 5.964 7.368
FLEx (Ours) 6.360 7.993
Figure 2: The figure on the left visualizes the distribution and central tendency of Helpfulness and Harmlessness scores for different methods (Original Model, FLEx, Local). Scatter points and kernel density contours represent the data distribution, with larger markers indicating the average scores. The table on the right presents the average Helpfulness and Harmlessness scores across evaluated methods, demonstrating performance in these two key aspects of open-ended generation. FLEx (Ours) achieves the highest average scores in both Helpfulness (6.360) and Harmlessness (7.993), showing significant improvement, particularly in harmlessness.
In this section, we evaluate the performance of FLEx on an open-ended benchmark Vicuna from helpfulness and harmlessness in two aspects. This experiment was conducted on a set of 20 clients under an IID (Independent and Identically Distributed) scenario to assess the performance of FLEx in open-ended generation tasks, specifically focusing on the aspects of Helpfulness and Harmlessness. The goal of this experiment was to evaluate the model’s ability to maintain high-quality, useful, and safe responses in a more balanced and homogeneous data setting. This IID setting allows for a more straightforward comparison of the different methods under conditions where the data distribution is uniform across clients, ensuring that the performance differences observed are primarily due to the methods themselves rather than data heterogeneity.
The left side of Figure 2 visualizes the distribution of Helpfulness and Harmlessness scores for Original Model, FLEx and Local. The scatter points represent individual evaluations, while the kernel density contours show the overall distribution of scores. The larger markers indicate the average scores for each method. We can see that FLEx not only achieves higher average scores but also exhibits a more concentrated and consistent performance, especially in the Harmlessness dimension. From the table on the right, we observe that FLEx outperforms all other methods in both Helpfulness (6.360) and Harmlessness (7.993), achieving the highest scores in both categories. This is particularly notable in the Harmlessness aspect, where FLEx significantly exceeds all other methods, including MoE + FedAdagrad (the second-best method).
These results demonstrate the advantages of FLEx in open-ended generation tasks, particularly in capturing the subtle distinctions of helpfulness and harmlessness. Compared to other methods, FLEx, through its personalized client models, is able to more finely capture the nuanced features of helpful and harmless content. In terms of helpfulness, FLEx performs exceptionally well, generating content that is not only valuable to users but also accurate and effective. In terms of harmlessness, FLEx stands out especially, as it effectively avoids generating potentially harmful or unsafe outputs, significantly surpassing other methods.
5.5 Communication Cost Analysis
We compare the communication cost of FLEx on MoE-based LLMs. Specifically, we examine the communication cost when applying the LoRA adapter to all MoE experts and transmitting all the experts across clients on Qwen1.5-MoE-A2.7B. The traditional FL cost is approximately 3.41% of the total parameters (i.e., 889M), while the FLEx cost is nearly 0.0659% (i.e., 17.2M) per communication round. It is evident that FLEx uses significantly lower communication costs to achieve high performance.
5.6 Ablation Study
Table 4: Ablation Study Results. ✓ indicates the component is included, and ✗ indicates it is excluded. Here, "No Personalized" denotes that each client prunes the expert using the C4 dataset, a pre-training dataset that does not contain personalized data. Conversely, "Personalized" refers to pruning the expert using the client’s own data.
Databricks-dolly-15k
Pruning Personalized Adaptive Gate Rouge-L
41.17
5.16
7.06
42.67
43.14
In this ablation study, we evaluate the contributions of the three core components of the FLEx framework: Pruning, Personalized Pruning, and Adaptive Gate. Experiments were conducted using the Databricks-dolly-15k dataset with FedAvg as the aggregation method and performance evaluated using the ROUGE-L metric. Results are summarized in Table 4. The baseline model, without any components enabled, achieved a ROUGE-L score of 41.17. Notably, enabling only Pruning drastically reduced the performance to 5.16, underscoring the critical necessity of personalization. Incorporating Personalized Pruning alongside Pruning slightly improved the score to 7.06, yet the performance remained limited, demonstrating that personalization alone is insufficient without effectively leveraging client-specific expertise. Enabling the Adaptive Gate raised performance substantially to 42.67, surpassing the baseline. Finally, integrating all three components—Pruning, Personalized Pruning, and Adaptive Gate yielded the highest performance of 43.14. This result underscores the critical role personalization plays within the FLEx framework, highlighting that its full potential is best realized when combined with the dynamic selection capability of adaptive gating to optimally accommodate diverse client datasets.
6 Conclusion
In this paper, we introduced FLEx, a federated learning framework tailored for MoE-based large language models. FLEx reduces communication and computation costs by allowing each client to fine-tune and retain a single, local expert while aggregating only the necessary updates from multi-head attention layers. This design effectively balances local personalization with global knowledge sharing, achieving up to 50$\times$ reduction in communication overhead compared to traditional methods. Extensive evaluations show that FLEx outperforms existing federated learning baselines on non-IID tasks, demonstrating its potential for efficient and scalable deployment of MoE-based LLMs in privacy-sensitive settings. Future work will focus on improving scalability and robustness across diverse applications.
Acknowledgement
This work was supported by National Natural Science Foundation of China (62303319, 62406195), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, Core Facility Platform of Computer Science and Communication of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
Supplementary organization:
Appendix A Algorithm Framework
We now present a detailed overview of our FLEx (Federated LLMs with Personalized Experts) algorithm framework. As illustrated in Algorithm 1, we incorporate personalized expert selection, execute local model updates for each client, and aggregate the attention-based updates at the server level. This process effectively balances global aggregation and local personalization. Algorithm 1 outlines the FLEx framework, beginning with personalized expert selection for each client. Each client utilizes their respective local dataset to perform expert selection, subsequently updating their local model using a personalized MoE layer. During global aggregation phases, clients transmit only their attention layer LoRA adapters to the central server, retaining personalized MoE layers locally. These adapters are aggregated across all clients, and the resulting global attention layer is integrated locally within each personalized MoE model.
Algorithm 1 FLEx: Federated LLMs with Personalized Experts
1:Input: Pre-trained MoE model, client set $\mathcal{C}$
2:for each client $i\in\mathcal{C}$ do
3:     Perform personalized expert selection (Algorithm 2)
4:     Update local model with personalized MoE layer
5:end for
6:for each global iteration do
7:     for each client $i\in\mathcal{C}$ do
8:         Send attention layer LoRA adapters to server
9:         Keep personalized MoE layer locally
10:     end for
11:     Aggregate global attention LoRA adapters
12:     Send global attention to clients and integrate into personalized models locally
13:end for
14:Output: Federated model with personalized experts on each client
Algorithm 2 provides the detailed procedure for personalized expert selection. Specifically, we employ a reconstruction loss strategy to prune the original MoE layers down to a tailored subset of experts uniquely optimized for each client. The selected subset effectively minimizes the discrepancy between the outputs of the full expert set and those produced by the personalized subset, thus ensuring optimal local performance.
Algorithm 2 Personalized Expert Selection
1:Input: Local dataset $\mathcal{D}_{i}$ for client $i$, pre-trained MoE weights, number of experts $K$, size of pruned subset $n$
2:Initialize local model with pre-trained MoE weights
3:for each MoE layer independently do
4:     for each input $x\in\mathcal{D}_{i}$ do
5:         Compute the outputs of all experts: $F(x)$
6:     end for
7:     Select personalized expert subset $S_{i}$ by solving:
$S_{i}=\arg\min_{S\subset\{1,\dots,K\},|S|=n}\left(\frac{1}{|\mathcal{D}_{i}|}\sum_{x\in\mathcal{D}_{i}}\|F_{S}(x)-F(x)\|_{F}\right)$
8:     Prune the MoE layer by retaining only experts in $S_{i}$
9:end for
10:Output: Personalized expert subsets and updated MoE layers
Appendix B Discussion
Our method retains the pre-trained weights in the MoE layer while adding a personalized side expert. This design maximally preserves pre-trained knowledge, leverages personalized experts to enhance client-specific capabilities, benefits from shared global knowledge aggregation, and significantly reduces communication overhead.
Limitation.
Our method adopts a layer-wise greedy pruning strategy, which ensures minimal reconstruction loss at each individual layer. However, this localized optimization does not necessarily guarantee a globally optimal pruning configuration across the entire model. As a result, the pruning decisions made at early layers may limit the potential of subsequent layers, leading to suboptimal global performance. Exploring more globally-aware pruning strategies across layers that better balance efficiency and model fidelity remains an important and underexplored direction for future research.
Further Comparsion with Previous Works.
It is worth mentioning that [15] proposed a federated learning approach specifically targeting frameworks utilizing Switch Transformers to address data heterogeneity. However, their approach involves fine-tuning all MoE layers, inevitably leading to significant forgetting from the pretrained models. This forgetting issue is particularly concerning given that current decoder-only large language models are pretrained on billions of tokens, making the preservation of pretrained knowledge an important task in federated learning. In contrast, our proposed method maintains the integrity of the original pretrained model parameters by introducing a specialized side-expert specifically designed to handle data heterogeneity. Our approach effectively mitigates the forgetting issue inherent in methods that fine-tune all MoE layers. Furthermore, we validate the applicability and efficacy of our federated learning approach through extensive experiments on billion-token pretrained, instruction-tuned large language models, including Qwen and DeepSeek.
Appendix C Further Experiments
C.1 Evaluation Across Different Model Architectures
In this section, we validate the robustness of FLEx across different model architectures. Specifically, we conduct experiments using the Databricks-dolly-15k dataset in combination with an alternative MoE-based model, DeepSeek-MoE-16B-Base [^3] https://huggingface.co/deepseek-ai/deepseek-moe-16b-base. We assess performance under a pathological non-IID scenario, with the results summarized in Table 5. From the table, we observe significant improvements across most evaluated tasks, with only a slight reduction in classification performance. When considering the average performance gains, our proposed method demonstrates a relatively substantial improvement (32.66) compared to the closest competitor, MoE+FedYogi (30.85). These results confirm that our approach maintains effectiveness and consistently delivers enhanced performance across varying model architectures.
Table 5: Experiments under another MoE-based LLM, DeepSeek-MoE-16B-Base. The following table shown Rouge-L performance comparison of different federated learning strategies on the Databricks-dolly-15k dataset under a pathological non-IID scenario. The best result over federated method is bolded and the second-best result is underlined. CLF: Classification, CQA: Closed Question Answering, IE: Information Extraction, Summ: Summarization
Databricks-dolly-15k
Method CLF CQA IE Summ Avg
Local Training 51.71 14.17 17.70 37.15 30.18
MoE + FedAvg 48.33 14.74 17.89 31.92 28.97
MoE + FedAvgM 48.25 14.21 17.99 32.55 28.25
MoE + FedAdam 49.96 19.15 20.30 33.83 30.81
MoE + FedAdagrad 50.10 18.75 20.83 32.91 30.64
MoE + SCAFFOLD 49.35 14.34 17.41 31.30 28.10
MoE + FedProx 49.35 14.55 17.50 32.17 28.39
MoE + FedYogi 50.63 18.42 20.70 33.66 30.85
FLEx (Ours) 50.33 20.11 21.29 38.93 32.66
C.2 Experiments on Trainable Modules
To further validate the effectiveness of FLEx’s training and sharing strategies, we conducted additional experiments investigating various combinations of trainable modules. Specifically, we explored the following approaches:
  1. 1.
    We select a single expert randomly at each layer for keeping same trainable parameters with FLEx.
  2. 2.
    We select all experts at each layer instead of choosing a single pruned expert, which consequently leads to higher communication costs and an increased number of trainable parameters.
The experimental results, summarized in Table 6, demonstrate that our proposed method achieves comparable performance while significantly reducing both communication and computational costs.
Table 6: Rouge-L performance comparison of different federated learning strategies, including our proposed method, on the Databricks-dolly-15k dataset with Qwen1.5-MoE-A2.7B under a pathological non-IID scenario. The best result over federated method is bolded and the second-best result is underlined. CP: Communication Parameter, TP: Trainable Parameter. denote that further shared and averaging all LoRA layers on MoE layers, which lead to more communication costs during training.
Databricks-dolly-15k
Method Modules CLF CQA IE Summ Avg CP(%$\downarrow$ TP(%$\downarrow$
MoE + FedAvg Single Expert 50.83 34.34 36.26 41.59 40.75 0.1327 0.1327
MoE + FedAvg All Experts 51.80 35.91 38.91 42.12 42.18 3.4172 3.4172
MoE + FedAvg Single Expert 51.48 34.77 37.23 41.74 41.30 0.0672 0.1327
MoE + FedAvg All Experts 51.10 32.51 41.39 40.60 41.40 0.0672 3.4172
FLEx (Ours) Single Experts 52.10 38.23 40.13 42.09 43.13 0.0672 0.1307
C.3 Comparison with Traditional Federated Learning
Additionally, we compare our method with traditional federated learning algorithms that treat MoE as dense models from two perspectives: performance and communication cost. The experiment is conducted using the Qwen1.5-MoE-A2.7B model on the Databricks-dolly-15k dataset under a challenging pathological non-IID scenario. As illustrated in Figure 3, our approach achieves comparable performance with significantly lower communication overhead.
Figure 3: Our method significantly reduces communication overhead while simultaneously enhancing performance, a stark contrast to traditional federated learning approaches that treat MoE as dense models, leading to prohibitive communication costs.
C.4 Experiments on Expert Aggregation Strategies
In this section, we evaluate several expert aggregation strategies within FLEx. In this paper, FLEx employs a sigmoid function as presented in formula (8). We have experimented with various aggregation strategies for side experts, comparing the sigmoid function against alternative approaches, such as ReLU and Tanh functions. Specifically, the ReLU-based weighting for side expert $g_{e}^{t}$ can be expressed as:
$g_{e}^{t}$ $=\text{ReLU}(\text{router}_{e}(u_{t}^{l})).$ (9)
Similarly, for the Tanh-based aggregation, the weight $g_{e}^{t}$ is calculated as:
$g_{e}^{t}$ $=\text{Tanh}(\text{router}_{e}(u_{t}^{l})).$ (10)
As shown in Table 7, our sigmoid-based aggregation method consistently demonstrates superior effectiveness. In contrast, the use of ReLU and Tanh functions results in suboptimal performance across aggregation strategies. From this result, we discover that the aggregation strategy of side expert requires carefully chosen.
Table 7: Experiments under different side expert aggregation strategies. Rouge-L performance on the Databricks-dolly-15k dataset with Qwen1.5-MoE-A2.7B under pathological non-IID scenario.
Databricks-dolly-15k
Method CLF CQA IE Summ Avg
ReLU 51.16 33.02 37.51 41.78 40.86
Tanh 41.40 20.40 31.34 37.88 32.75
Sigmoid 52.10 38.23 40.13 42.09 43.13
Figure 4: Activation counts of experts in the Qwen1.5-MoE-A2.7B model evaluated on the C4 dataset.
C.5 Experiments on Expert Load Balancing
We conducted experiments to evaluate the activation frequency of experts within the Qwen1.5-MoE-A2.7B model using the C4 dataset [^4] https://huggingface.co/datasets/allenai/c4 across different methods. Figure 4 illustrates the activation count for individual experts. This analysis demonstrates that our method achieves a more balanced workload distribution across different experts. Specifically, the original model and the FedAvg method tend to produce uneven workloads, causing certain experts, such as expert 31 and expert 43, to become significantly overloaded. Specifically, the original model exhibits a standard deviation of 1859.80, while FedAvg demonstrates a slightly higher imbalance with a standard deviation of 1906.85. In contrast, our proposed FLEx method significantly improves load balancing among experts, achieving a notably lower standard deviation of 1259.88.
C.6 Data Distribution for Non-IID
In this paper, we explore different scenarios for data distribution, specifically focusing on pathological non-IID and Dirichlet non-IID settings. As illustrated in Figure 5, the pathological non-IID scenario assigns four distinct tasks to four separate clients, resulting in an extreme non-IID environment. This scenario significantly complicates the aggregation of models across clients. The Dirichlet non-IID distribution offers a compromise between the pathological non-IID and IID distributions. Specifically, a lower value of the hyperparameter $\alpha$ indicates a higher degree of non-IID distribution among clients. Unless explicitly stated otherwise, this paper defaults to the pathological non-IID distribution.
Figure 5: Distribution of instruction types across different clients in federated learning under various settings. The first panel shows the pathological distribution, where each client has a highly imbalanced data distribution. The second panel illustrates the Dirichlet distribution with alpha=0.1, resulting in a more skewed distribution across clients. The third panel presents the Dirichlet distribution with alpha=1.0, demonstrating a more uniform distribution of categories (classification, closed question answering, information extraction, and summarization) across clients.
C.7 Overview of Hyperparameters
In this section, we present the hyperparameters used for the models Qwen1.5-MoE-A2.7B and DeepSeek-MoE-16B-Base. Detailed hyperparameter configurations are summarized in Table 8.
Model Name Qwen1.5-MoE-A2.7B DeepSeek-MoE-16B-Base
# Params 14.3B 16.4B
# Layers 24 28
Hidden Size 2048 2048
# Attn Heads 16 16
# Shared Experts 1 2
# Routed Experts 60 (4 activated) 64 (6 activated)
Relative Expert Size 0.25 0.25
Sequence Length 2048 2048
Learning Rate 4e-5 4e-5
Table 8: Model hyperparameter configurations.
C.8 Dataset Information
Databricks-dolly-15k.
The Databricks-dolly-15k dataset is an open-source collection consisting of 15k instruction-following examples generated collaboratively by Databricks employees. In our analysis, we specifically utilize data from the classification, closed QA, information extraction, and summarization categories.
Alpaca-gpt4.
The Alpaca-gpt4 dataset comprises 52k English-language instruction-following instances generated by GPT-4, based on prompts originally crafted for the Alpaca dataset. This dataset mirrors the structure of the Alpaca data but leverages GPT-4’s advanced generative capabilities for creating more nuanced outputs suitable for fine-tuning large language models.
Finance-Alpaca.
The Finance-Alpaca dataset merges Stanford’s Alpaca dataset and the FiQA financial question-answering dataset, supplemented with an additional 1,300 instruction-response pairs uniquely generated using GPT-3.5. In total, it encompasses 68,912 QA pairs designed specifically for applications in finance-related natural language processing tasks.
MedAlpaca.
The MedAlpaca dataset provides comprehensive coverage of medical domains, incorporating fundamental medical sciences, clinical knowledge, and clinical reasoning skills. It contains a total of 33,955 QA pairs designed to support the training and evaluation of medical-domain-specific language models and AI applications.
Vicuna Benchmark.
The Vicuna benchmark employs an innovative ’LLM-as-a-judge’ methodology, presenting 80 carefully curated questions spanning multiple categories, including general knowledge, coding challenges, role-playing, creative writing, common sense reasoning, and complex problem-solving. This approach enables a multidimensional evaluation of large language models, highlighting their capabilities across diverse cognitive tasks.
Appendix D Prompt Template
D.1 Alpaca Template
We utilize the Alpaca template as the standardized format for our instruction tuning, which helps define the structure of prompts and responses for more efficient model training.
Below is an instruction that describes a task. Write a response that appropriately completes the request.
### Instruction:
### Response:
D.2 Vicuna Helpful Template
To effectively evaluate helpfulness in open-ended tests, we use this template to configure powerful models as judges.
[Instruction]
Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 (You score can be a float with two decimal place) by strictly following this format: "[[rating]]", for example: "Rating: [[5.62]]".
[Question]
{question}
[The Start of Assistants Answer]
{answer}
[The End of Assistants Answer]
D.3 Vicuna Harmless Template
For open-ended testing, this template guides powerful models acting as judges in assessing harmlessness.
[Instruction]
You are an impartial judge evaluating whether an AI assistant s response contains harmful content. Be extremely strict even minor issues should lower the score. Your evaluation should consider factors such as inaccuracy, impoliteness, subjectivity, irrelevance, offensiveness, sensitivity, harmfulness, uselessness, or misleading nature. Begin your evaluation by providing a short explanation. Be as strict as possible. After providing your explanation, you must rate the response on a scale of 1 (highly dangerous) to 10 (harmless). (Your score can be a float with two decimal places) by strictly following this format: "[[rating]]", for example: "Rating: [[5.62]]".
[Question]
{question}
[The Start of Assistants Answer]
{answer}
[The End of Assistants Answer]